### v1

# UniADC: A Unified Framework for Anomaly Detection and Classification

Abstract
In this paper, we introduce the task of unified anomaly detection and classification, which aims to simultaneously detect anomalous regions in images and identify their specific categories. Existing methods typically treat anomaly detection and classification as separate tasks, thereby neglecting their inherent correlation, limiting information sharing, and resulting in suboptimal performance. To address this, we propose UniADC, a Unified Anomaly Detection and Classification model that can effectively perform both tasks with only a few or even no anomaly images. Specifically, UniADC consists of two key components: a training-free controllable inpainting network and a multi-task discriminator. The inpainting network can synthesize anomaly images of specific categories by repainting normal regions guided by anomaly priors, and can also repaint few-shot anomaly samples to augment the available anomaly data. The multi-task discriminator is then trained on these synthesized samples, enabling precise anomaly detection and classification by aligning fine-grained image features with anomaly-category embeddings. We conduct extensive experiments on three anomaly detection and classification datasets, including MVTec-FS, MTD, and WFDD, and the results demonstrate that UniADC consistently outperforms existing methods in anomaly detection, localization, and classification. The code is available at https://github.com/cnulab/UniADC.
1 Introduction
Figure 1: Task settings for unified anomaly detection and classification. (a) Zero-shot: training with a few normal samples and prior knowledge for each anomaly category. (b) Few-shot: training with a few normal samples and few-shot anomaly samples per category. (c) Inference: UniADC predicts anomaly scores and categories for query images.
Image anomaly detection aims to train models to detect and localize anomalous regions within images. It has numerous practical applications and has attracted growing research interest in recent years [37, 40, 5, 22]. Anomaly classification further categorizes detected anomalies, which aids in assessing anomaly severity and facilitating root-cause analysis. Existing studies [23, 13] typically treat anomaly classification as a downstream task of anomaly detection, in which the detected anomalous regions are cropped into sub-images and then fed into a separate classification model. This two-stage pipeline ignores the inherent connection between anomaly detection and classification, and hinders information sharing across the two tasks. Moreover, missed and over detections during the detection stage, as well as the difficulty in selecting appropriate anomaly score thresholds, pose a series of challenges for subsequent classification. As a result, such methods often suffer from high implementation complexity and suboptimal performance, limiting their practicality in real-world scenarios.
In this paper, we integrate anomaly detection and classification tasks and address them with a unified model. We propose zero-shot and few-shot settings for unified anomaly detection and classification, both extending the widely studied few-shot anomaly detection task [14, 20, 8, 36]. Few-shot anomaly detection provides only a small number of normal samples for model training, thereby reducing data collection costs, and has been widely applied in industrial [14, 20, 8] and medical [45, 11] fields. This setting is highly aligned with anomaly classification tasks in which available data is limited. Building upon this, the zero-shot anomaly detection and classification task does not provide additional anomaly samples for model training, but instead offers prior knowledge about anomaly categories, such as textual descriptions or shape information. This setting simulates real-world scenarios where collecting anomalous data is challenging. Differently, the few-shot anomaly detection and classification task provides a small number of anomaly samples for each category during model training, simulating practical scenarios where anomalous data is limited. Further explanation of the two settings is provided in Fig. 1.
Figure 2: Comparison of UniADC with alternative methods on the MVTec-FS dataset. UniADC demonstrates superior performance over existing methods in Anomaly Detection (AD), Anomaly Localization (AL), Anomaly Classification (AC), and Pixel-level Anomaly Classification (PAC).
To address the above tasks, we propose UniADC, a unified model for anomaly detection and classification, which tackles the challenge of scarce anomalous data through controllable image inpainting. It consists of two key components: a training-free controllable inpainting network and a multi-task discriminator. The controllable inpainting network consists of a pre-trained latent diffusion model [26] and an inpainting control network [44], enabling controllable inpainting guided by either anomaly priors or anomaly samples. Specifically, anomaly prior-guided controllable inpainting synthesizes category-specific anomalous images by repainting normal regions according to the provided prior knowledge of anomalies, whereas anomaly sample-guided controllable inpainting refines and repaints few-shot anomalous samples to enhance data diversity. In addition, we propose a category consistency selection strategy to filter synthetic anomaly samples that are highly consistent with the target category. The multi-task discriminator is trained on these synthetic anomaly samples and leverages the powerful generalization capability of the existing vision-language models such as CLIP [25] or dino.txt [16, 30] to achieve end-to-end anomaly detection and classification. As shown in Fig. 2, UniADC comprehensively outperforms existing methods in anomaly detection, localization, and classification, particularly strengthening the reliability of defect categorization. Our main contributions are summarized as follows:
  • We introduce the task of unified anomaly detection and classification, which has broad application prospects yet remains underexplored. To this end, we propose UniADC, which enables accurate anomaly detection and classification under both zero-shot and few-shot settings.
  • We propose a controllable inpainting network that can generate category-specific anomaly samples conditioned on either anomaly prior or few-shot anomaly images, without additional training. This enables broad applicability across various anomaly detection and classification tasks, serving as an effective alternative to existing anomaly synthesis methods.
  • Extensive experiments on the MVTec-FS [23], MTD [12], and WFDD [3] demonstrate the effectiveness of UniADC and highlight its potential for real-world anomaly detection and classification applications.
Figure 3: Overview of the proposed UniADC pipeline, which consists of a training-free controllable inpainting network and a multi-task discriminator. The controllable inpainting network supports two modes: anomaly prior-guided controllable inpainting and anomaly sample-guided controllable inpainting, enabling the generation of category-specific anomaly samples under different settings. The multi-task discriminator aligns fine-grained image features with anomaly category embeddings for accurate anomaly detection and classification.
2 Related Work
Few-shot Anomaly Detection. Few-shot anomaly detection methods train models using only a small number of normal samples and often rely on the general knowledge of foundation models to overcome the challenges posed by limited data. WinCLIP [14] employs the CLIP model [25] and aligns image patches with normal and anomalous prompts via sliding windows, enabling anomaly detection and fine-grained anomaly localization. PromptAD [20] introduces prompt learning to enhance the alignment process and refine decision boundaries. AnomalyGPT [8] feeds the query images and their anomaly scores into an LLM for anomaly inspection, eliminating the dependency on decision thresholds. It is worth noting that some existing methods also incorporate certain prior knowledge of anomalies to improve anomaly detection. For example, PromptAD [20] employs textual descriptions of each anomaly category to improve detection accuracy. AnomalyGPT [8] predefines the expected appearance of query images and potential anomaly types. AnoGen [9] constrains the size range of synthesized anomalies according to their types. In contrast, the proposed UniADC leverages anomaly prior knowledge for both anomaly detection and classification, providing a functional extension to existing methods.
Anomaly Classification. Anomaly classification aims to identify the categories of detected anomalous regions, which remains highly challenging due to the scarcity of anomaly samples. ZipAF [23] employs the AlphaCLIP model [35] to extract region-contextual anomaly features and adopts a zero-initialized projection to align query features with cached anomaly features for few-shot anomaly classification. However, it does not incorporate the anomaly detection process, which may result in degraded performance in practical scenarios. AnomalyNCD [13] integrates novel class discovery methods into the existing anomaly detection pipeline for anomaly detection and clustering. Recent work [28] proposes a zero-shot multi-type anomaly detection method, MultiADS, which trains the model to learn cross-domain anomaly priors for anomaly classification. However, due to the lack of targeted training, this method struggles to identify domain-specific anomaly types, such as identifying highly similar defects within a specific product [33, 12, 43].
Anomaly Synthesis. Existing anomaly synthesis methods can generally be classified into zero-shot and few-shot approaches, depending on the availability of real anomaly data during the synthesis process [18]. Zero-shot approaches generate anomaly samples relying on predefined data augmentation rules [19, 29], noise injection [38, 46], or textual descriptions [34], without access to real anomaly samples. However, the synthesized anomalies may exhibit a distribution shift from real-world cases, limiting their effectiveness. In contrast, few-shot approaches [7, 10, 15, 32] aim to enhance data diversity by augmenting a limited number of real anomaly samples. To the best of our knowledge, UniADC is the first method capable of synthesizing anomalies under both zero-shot and few-shot settings, thereby significantly enhancing its applicability across a wide range of anomaly detection and classification tasks.
3 Method
In this section, we present our proposed anomaly detection and classification model, UniADC, with its overall pipeline illustrated in Fig. 3. We first provide the problem definition, then describe the main functionalities and implementation details of the controllable inpainting network, including anomaly prior-guided controllable inpainting and anomaly sample-guided controllable inpainting. Finally, we introduce the multi-task discriminator of UniADC.
3.1 Problem Definition
In few-shot anomaly detection, the model is provided with a support set of normal samples $\mathcal{D}^{n}=\{X_{1},X_{2},\cdots,X_{K_{n}}\}$, where $K_{n}$ denotes the number of normal samples per image class. Given a query image, the model predicts an image-level anomaly detection score $I_{d}\in[0,1]$ and a pixel-level anomaly map $S_{d}\in[0,1]^{H\times W}$, where $H$ and $W$ denote the image height and width, respectively. The zero-shot and few-shot anomaly detection and classification tasks respectively provide additional anomaly prior knowledge and an anomaly support set for model training. Specifically, the anomaly support set is represented as $\mathcal{D}^{a}=\bigcup_{y\in\{1,2,\cdots,Y\}}{\{(X_{1}^{y},M_{1}^{y}),\>(X_{2}^{y},M_{2}^{y}),\cdots,\>(X_{K_{a}}^{y},M_{K_{a}}^{y})\}}$, where $Y$ denotes the number of anomaly categories, $K_{a}$ is the number of samples per anomaly category, and $M$ denotes the binary anomaly mask. The model is further trained to predict both an image-level classification result $I_{c}\in\{0,1,\cdots,Y\}$ and a pixel-level classification result $S_{c}\in\{0,1,\cdots,Y\}^{H\times W}$, where 0 indicates the normal class, and 1 to $Y$ represent the $Y$ anomaly categories.
3.2 Anomaly Prior-Guided Controllable Inpainting
This module synthesizes category-specific anomaly samples based on provided anomaly prior, making it suitable for zero-shot anomaly detection and classification tasks. The anomaly prior includes the shape, size, and textual description of each anomaly category. Shape and size are used to generate the anomaly mask, while the textual description serves as a prompt to guide the synthesis of the desired anomalous appearance. Specifically, we define eight types of candidate masks to simulate various anomaly shapes, including Rectangle, Line, Polygon, Ellipse, Hollow Ellipse, Random Brush, Perlin Noise, and Foreground Mask. Except for Foreground Mask, each type is generated at three scales: Large, Medium, and Small. We use specific mask types for mask generation based on the given anomaly shape and size. For example, we specify the shape of the anomaly category “Hole” as Ellipse, which will generate elliptical masks of arbitrary sizes. When both shape and size priors are unavailable for an anomaly category, all mask types are used for mask generation.
For a given normal image $X$ and prior knowledge from anomaly category $y$, we first generate an anomaly mask $M^{y}$ under the shape-size constraints. Then, the latent diffusion model encodes the image $X$ and the text prompt $p^{y}$ into latent variables $z=\mathbb{E}_{v}(X)$ and $z^{p}=\mathbb{E}_{t}(p^{y})$, where $\mathbb{E}_{v}(\cdot)$ and $\mathbb{E}_{t}(\cdot)$ denote the vision and text encoders, respectively. We define the number of forward diffusion steps as $T^{\prime}=T\cdot\gamma$, where $T$ denotes the total number of diffusion steps in the original latent diffusion model, and $\gamma\in(0,1]$ controls the noise strength. A larger $\gamma$ injects more noise into the latent variable, resulting in synthesized anomalies that deviate further from the original image distribution and exhibit more visually prominent anomalous patterns. To ensure spatial consistency between synthesized anomalous and the anomaly mask, we use an inpainting control network to precisely control the locations of the synthesized anomalies. Specifically, at each time step $t\in\{1,2,\cdots,T^{\prime}\}$, the inpainting control network $\phi$ predicts the conditioning variable $z^{m}=\phi(z_{t},t,z,M^{y})$ of noisy latent $z_{t}$ for controlling the denoising process:
$z_{t-1}=\Psi(z_{t},t,z^{p},z^{m})$ (1)
where $\Psi$ denotes the latent diffusion model, and the conditioning variables $z^{p}$ and $z^{m}$ respectively control the semantics and spatial positions of the synthesized anomalies. The inpainting control network ensures that the denoised latent representation $z_{0}$ remains consistent with the original latent $z$ outside the masked region, while repainting only the masked region to align with the prompt embedding $z^{p}$. Finally, the denoised latent representation $z_{0}$ is decoded by decoder $\mathbb{D}(\cdot)$ into the anomaly image $X^{y}=\mathbb{D}(z_{0})$. In our implementation, we adopt Stable Diffusion v1.5 [26] as the latent diffusion model and BrushNet [17] as the inpainting control network. Both models are trained on high-quality and diverse data, enabling UniADC to generalize effectively across diverse anomaly synthesis scenarios.
However, the above synthesis process does not ensure the semantic alignment between anomaly masks and their corresponding textual descriptions, which may lead to inconsistencies between the synthesized anomaly samples and their intended categories. For example, in the case of “a broken transistor lead”, the synthesized anomaly should be located in the lead region, rather than randomly elsewhere. To this end, we propose a category consistency selection strategy to further filter synthetic anomalous images with high category consistency, thereby eliminating noisy samples introduced by anomaly synthesis. Given an anomaly category $y$, we first generate a mini-batch of synthesized samples using diverse anomaly masks, denoted as $\mathcal{S}^{y}=\{(X_{b}^{y},M_{b}^{y})\}_{b=1}^{B}$, where $B$ is the mini-batch size. We then adopt the AlphaCLIP model [35] to evaluate the semantic consistency between the synthesized image-mask pairs and the textual description of the anomaly category. For the synthetic anomaly sample $(X^{y}_{b},M^{y}_{b})$, we calculate its category matching score as:
[…]
y{1,2,,Y}
[…]
where $\psi_{v}$ and $\psi_{t}$ denote the vision and text encoders of AlphaCLIP, respectively, and $\langle\cdot,\cdot\rangle$ denotes the inner product. The AlphaCLIP model aligns the image region features guided by the masks with the text embeddings. If the synthetic anomalous region fails to match the corresponding anomaly description, such as being placed in the wrong location, the matching score decreases. In the end, we select the synthetic anomaly sample with the highest matching score in the mini-batch $\mathcal{S}^{y}$ for discriminator training.
Table 1: Comparison of UniADC with alternative methods under zero-shot anomaly detection and classification settings. $\mathcal{y}$ denotes methods that use anomaly priors.
$K_{n}\;\;$ Method MVTec-FS [23] MTD [12] WFDD [3]
I-AUC P-AUC PRO Acc mIoU I-AUC P-AUC PRO Acc mIoU I-AUC P-AUC PRO Acc mIoU
1 InCTRL [49] 92.36 - - - - 70.34 - - - - 96.66 - - - -
PatchCore [27] 84.76 93.77 82.44 - - 68.22 73.44 59.50 - - 84.81 96.07 70.62 - -
$\mathcal{y}$ PromptAD [20] 91.80 95.09 87.06 - - 86.27 71.70 70.61 - - 96.90 97.10 86.72 - -
$\mathcal{y}$ AnomalyGPT [8] 93.48 95.92 86.60 45.85 - 71.91 68.41 58.67 27.57 - 96.89 97.72 85.69 45.76 -
$\mathcal{y}$ WinCLIP [14] 93.20 94.43 86.47 40.75 25.17 77.53 69.26 57.58 28.69 15.10 95.72 93.94 78.98 34.64 27.53
$\mathcal{y}$ UniADC(CLIP) 95.03 96.33 88.75 66.07 32.90 86.19 80.80 77.69 55.82 27.92 97.01 99.00 87.03 86.70 45.55
$\mathcal{y}$ UniADC(DINO) 96.37 96.11 89.16 68.30 35.06 90.09 82.65 79.98 59.22 29.58 98.08 99.46 92.13 88.88 52.17
2 InCTRL [49] 93.01 - - - - 72.07 - - - - 97.36 - - - -
PatchCore [27] 88.49 94.43 84.65 - - 69.90 75.07 60.52 - - 88.06 96.35 71.45 - -
$\mathcal{y}$ PromptAD [20] 93.95 95.42 87.93 - - 87.06 73.53 69.91 - - 97.15 97.20 86.81 - -
$\mathcal{y}$ AnomalyGPT [8] 94.91 96.24 87.97 52.65 - 72.57 70.20 60.12 32.56 - 97.48 97.78 85.75 47.39 -
$\mathcal{y}$ WinCLIP [14] 94.37 94.60 86.95 41.93 25.34 78.07 71.57 57.70 29.90 15.28 96.50 94.16 80.19 35.83 27.62
$\mathcal{y}$ UniADC(CLIP) 95.27 96.53 88.52 71.64 35.12 90.31 82.92 78.99 62.87 30.56 97.55 98.98 87.14 89.48 48.15
$\mathcal{y}$ UniADC(DINO) 97.09 97.04 92.15 74.74 36.66 92.30 83.52 81.22 64.15 31.85 98.21 99.46 92.35 89.22 53.38
4 InCTRL [49] 93.62 - - - - 73.23 - - - - 97.38 - - - -
PatchCore [27] 90.75 95.28 86.56 - - 71.40 75.35 59.67 - - 88.26 97.21 71.41 - -
$\mathcal{y}$ PromptAD [20] 94.91 95.92 89.86 - - 88.16 73.82 71.49 - - 96.79 97.53 87.80 - -
$\mathcal{y}$ AnomalyGPT [8] 96.10 96.42 91.09 56.59 - 74.38 68.68 59.92 42.48 - 98.00 97.92 86.35 48.24 -
$\mathcal{y}$ WinCLIP [14] 95.17 94.98 87.67 42.69 25.70 78.90 69.84 58.11 29.91 16.17 96.74 94.58 80.28 34.99 27.79
$\mathcal{y}$ UniADC(CLIP) 96.18 96.47 91.17 73.33 35.14 90.88 84.88 79.81 63.98 31.29 98.15 99.09 88.29 90.85 48.24
$\mathcal{y}$ UniADC(DINO) 97.65 97.36 92.68 76.35 37.23 93.04 85.82 81.43 66.03 33.75 98.60 99.39 94.76 91.06 55.44
3.3 Anomaly Sample-Guided Controllable Inpainting
Anomaly sample-guided controllable inpainting aims to enrich the diversity of anomalies by repainting few-shot anomaly samples, making it well-suited for few-shot anomaly detection and classification tasks. Given a normal image $X$ and an anomalous sample $(X^{y},M^{y})$, we first crop the anomalous region from $X^{y}$ based on the mask $M^{y}$, and paste it at a random location on the normal image $X$, yielding a preliminary synthetic anomaly image $\hat{X}^{y}$ and its corresponding mask $\hat{M}^{y}$. To increase the shape variation, we apply identical data augmentations, such as affine transformations, to both the anomalous region and its mask. However, at this stage, the synthesized anomalies still suffer from limited diversity and suboptimal visual coherence. To improve diversity and image quality, we use the latent diffusion model along with an inpainting control network to repaint the pasted regions in $\hat{X}^{y}$. The repainting process is guided by simple text prompts (e.g., the name of the anomaly category) and follows the same diffusion and denoising procedures as described in anomaly prior-guided inpainting. The degree of diversity and deviation from the original anomaly is controlled by the noise factor $\gamma$. To ensure semantic consistency, we perform category consistency selection by computing the Structural Similarity Index (SSIM) [39] between the original and repainted anomalous sub-images. These sub-images include both the anomalous region and a fixed amount of surrounding normal context. We compute the SSIM-based category matching score for each sample and retain the highest-scoring synthetic anomaly sample for discriminator training.
Table 2: Comparison of UniADC with alternative methods under few-shot anomaly detection and classification settings.
$K_{n}$ $K_{a}$ Method MVTec-FS [23] MTD [12] WFDD [3]
I-AUC P-AUC PRO Acc mIoU I-AUC P-AUC PRO Acc mIoU I-AUC P-AUC PRO Acc mIoU
1 1 PatchCore+ZipAF [27] 84.76 93.77 82.44 52.10 37.30 68.22 73.44 59.50 28.26 20.33 84.81 96.07 70.62 61.75 39.10
PromptAD+ZipAF [20] 92.74 95.72 88.06 49.12 39.57 86.85 72.30 70.87 31.79 28.92 97.47 97.51 86.23 64.11 44.60
WinCLIP+ZipAF [14] 93.20 94.43 86.47 49.60 34.39 77.53 69.26 57.58 33.78 24.23 95.72 93.94 78.98 59.59 37.30
AnomalyCLIP+ZipAF [48] 95.59 96.00 87.98 52.60 40.25 76.18 75.98 69.54 25.61 21.78 93.76 98.59 87.58 76.48 45.99
UniADC(CLIP) 97.69 98.53 89.72 84.72 48.62 88.49 82.41 77.41 60.81 34.80 98.50 98.83 92.96 93.51 48.76
UniADC(DINO) 98.42 98.96 92.26 86.74 51.28 91.41 82.88 80.07 62.47 32.21 99.85 99.37 94.07 96.10 50.53
2 1 PatchCore+ZipAF [27] 88.49 94.43 84.65 56.16 39.63 69.90 75.07 60.52 30.15 21.03 88.06 96.35 71.45 63.82 39.71
PromptAD+ZipAF [20] 94.58 95.66 88.93 51.40 40.09 87.41 74.10 71.12 36.09 29.78 97.52 97.74 87.01 67.17 44.51
WinCLIP+ZipAF [14] 94.37 94.60 86.95 49.99 34.77 78.07 71.57 57.70 33.31 24.54 96.50 94.16 80.19 60.21 37.28
AnomalyCLIP+ZipAF [48] 95.94 96.12 88.77 53.80 40.90 76.42 76.11 70.62 26.57 21.17 94.96 98.75 87.80 77.99 46.49
UniADC(CLIP) 98.35 98.63 90.64 85.20 48.93 90.70 84.49 80.71 63.78 35.58 98.97 99.16 94.64 93.66 49.71
UniADC(DINO) 98.56 98.90 92.48 86.85 51.49 92.57 86.79 83.17 65.10 34.33 99.87 99.48 94.12 97.22 51.78
2 PatchCore+ZipAF [27] 88.49 94.43 84.65 58.25 40.32 69.90 75.07 60.52 34.71 21.78 88.06 96.35 71.45 64.52 41.29
PromptAD+ZipAF [20] 95.16 95.60 89.03 52.66 40.93 88.46 74.53 71.91 43.89 31.16 97.75 98.20 86.81 67.91 45.25
WinCLIP+ZipAF [14] 94.37 94.60 86.95 51.76 35.25 78.07 71.57 57.70 37.36 27.14 96.50 94.16 80.19 61.64 37.69
AnomalyCLIP+ZipAF [48] 95.93 96.47 89.12 52.93 42.49 76.72 76.32 73.90 35.57 24.81 95.25 98.73 90.38 81.17 47.74
UniADC(CLIP) 98.75 98.88 93.31 89.88 54.93 93.15 86.20 83.36 71.34 37.05 99.14 99.34 94.85 94.90 51.06
UniADC(DINO) 99.05 99.10 93.55 88.72 54.21 93.85 89.31 85.78 70.87 35.46 99.93 99.49 94.19 97.76 51.91
4 1 PatchCore+ZipAF [27] 90.75 95.28 86.56 60.32 39.80 71.40 75.35 59.67 31.58 21.92 88.26 97.21 71.41 64.17 42.93
PromptAD+ZipAF [20] 96.49 96.10 90.02 55.21 41.66 87.78 73.82 71.49 38.36 30.13 97.57 97.83 87.80 70.59 44.87
WinCLIP+ZipAF [14] 95.17 94.98 87.67 53.75 35.98 78.90 69.84 58.11 33.89 28.32 96.74 94.58 80.28 63.59 37.59
AnomalyCLIP+ZipAF [48] 95.98 96.67 89.77 55.07 40.83 76.84 76.22 71.11 29.58 22.37 95.22 98.74 90.17 78.88 46.97
UniADC(CLIP) 98.57 98.55 91.42 85.57 50.01 92.74 84.98 82.15 64.40 35.86 99.35 99.19 94.48 93.91 49.94
UniADC(DINO) 98.70 98.91 93.35 86.83 52.37 93.32 87.12 83.83 66.59 34.87 99.87 99.45 94.59 97.67 51.80
3.4 Multi-task Discriminator
We train the multi-task discriminator on synthetic anomalous samples so that it learns to align fine-grained image features with anomaly category embeddings, enabling unified anomaly detection and classification. First, we represent each synthetic anomalous sample as a 4-tuple $(X,y,M_{d},M_{c})$, where $X$ is the synthesized anomalous image, $y$ is the anomaly category label, and $M_{d}$ and $M_{c}$ respectively denote the anomaly detection mask and classification mask, for a pixel location $(i,j)$ satisfying:
[…]
We employ a pre-trained vision encoder to extract multi-scale image features $\{x_{1},x_{2},\cdots,x_{N}\}=\varphi(X)$, and fuse these features into a fine-grained representation through a feature fusion network $f=\theta(x_{1},x_{2},\cdots,x_{N})$, where $f\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times C}$. The fusion network follows a fully convolutional architecture that progressively fuses intermediate-layer features from high to low levels. We then employ a text encoder to embed the names of anomaly categories to obtain $\{g_{1},g_{2},\cdots,g_{Y}\}$, where each $g_{y}\in\mathbb{R}^{C}$ denotes the embedding of the $y$-th anomaly category. The similarity between the anomaly category embedding $g_{y}$ and the feature map $f$ at a spatial location $(h,w)$ is computed as:
$s_{y}(h,w)=\sigma(\langle f(h,w),g_{y}\rangle/\epsilon)$ (4)
where $\sigma(\cdot)$ denotes the Sigmoid function, and $\epsilon$ is a learnable scaling parameter. We compute the similarity score at each spatial location and upsample the resulting similarity matrix to the original image resolution, yielding a classification map $S_{y}\in(0,1)^{H\times W}$, which indicates the likelihood of anomaly category $y$ at each pixel. By computing classification maps for all categories, we obtain a set of classification maps $\{S_{1},S_{2},\cdots,S_{Y}\}$. The pixel-level anomaly detection map $S_{d}$ is then computed as the average of these classification maps: $S_{d}=\frac{1}{Y}\Sigma_{y=1}^{Y}S_{y}$. The maximum value of $S_{d}$ is used as the image-level anomaly detection score $I_{d}$. Then, the pixel-level anomaly classification result at a pixel location $(i,j)$ is:
Sc(i,j)={argmax
y{1,2,,Y}
Sy
(i,j)
,Sd(i,j)
τ
0,Sd(i,j)<τ
[…]
where $\tau$ represents the anomaly score threshold. We use the anomaly category with the largest pixel area in $S_{c}$ as the image-level anomaly classification result $I_{c}$. If all pixels in $S_{c}$ are classified as normal, then $I_{c}$ is set to 0. We optimize the feature fusion network using Binary Focal loss [21] and Dice loss [24] for anomaly detection, and Cross-Entropy loss for anomaly classification. The overall loss function is:
$\mathcal{L}$ $=\mathcal{L}_{Focal}(S_{d},M_{d})+\mathcal{L}_{Dice}(S_{d},M_{d})$ (6)
$+\lambda\mathcal{L}_{CE}([S_{1},S_{2},\cdots,S_{Y}],M_{c})$
where $\cdot$$\cdot$] denotes the concatenation operation, and $\lambda$ is a weight hyperparameter. The normal regions in $M_{c}$ are ignored during the computation of $\mathcal{L}_{CE}$. Unlike prior works [14, 20], we discard normal prompts to avoid performance degradation caused by the severe class imbalance between normal pixels and per-category anomalous pixels.
4 Experiments
4.1 Experimental Setup
Figure 4: Qualitative results of UniADC on the MVTec-FS dataset under $\text{zero-shot}_{\>(K_{n}=2)}$ and $\text{few-shot}_{\>(K_{n}=2,\,K_{a}=1)}$ settings.
Datasets. We conduct extensive evaluations on three anomaly detection and classification datasets, including MVTec-FS [23], MTD [12], and WFDD [3]. The MVTec-FS dataset is introduced by ZipAF [23] as an extension of the MVTec-AD benchmark [1] for anomaly classification. It contains 15 types of industrial products, each with an average of four anomaly categories. The MTD dataset [12] consists of 1,344 images of magnetic tiles, including five types of anomalies. The subtle differences among these anomalies make the classification task particularly challenging. The WFDD dataset [3] contains 4,101 images across four types of fabrics. We reclassified the dataset by grouping defects of the same type, resulting in an average of three anomaly types per fabric. More details of the WFDD dataset [3] are provided in the Appendix. Some popular anomaly detection datasets, such as VisA [50], are excluded from this work due to the lack of anomaly category labels.
Implementation Details. We adopt a consistent experimental setup for both anomaly prior-guided and anomaly sample-guided inpainting. Specifically, we use the DDIM sampler [31] with 1,000 original diffusion steps. The noise factor $\gamma$ is uniformly sampled between 0.4 and 0.6, and the number of accelerated sampling steps is set to 10. For mask generation, we leverage BiRefNet [47] for binary segmentation to ensure that the mask is located within the foreground region. We set the mini-batch size to 32 for category consistency selection, and generate 16 samples with a resolution of $512\times 512$ per anomaly category for discriminator training. For the multi-task discriminator, we adopt two experimental settings, namely UniADC(CLIP) and UniADC(DINO), which use the CLIP ViT-L/14 [25] and the DINOv3-based dino.txt [30] as the vision-language backbone of the discriminator, respectively. We set both the anomaly score threshold $\tau$ and the loss weight $\lambda$ to 0.5. The list of anomaly prior is provided in the Appendix.
Metrics. We use the Area Under the Receiver Operating Characteristic Curve (I-AUC) metric to evaluate image-level anomaly detection performance. For pixel-level anomaly location, we use Pixel-AUROC (P-AUC) and Per Region Overlap (PRO) [2] metrics. In addition, we report the Top-1 classification accuracy (Acc) and Mean Intersection over Union (mIoU) to evaluate image-level and pixel-level anomaly classification performance, respectively. For the “combined” category in MVTec-FS dataset [23], we only evaluate its anomaly detection performance and exclude it from the anomaly classification metrics, as in previous works [23, 13].
Baselines. In the zero-shot setting, we adopt InCTRL [49], PatchCore [27], PromptAD [20], AnomalyGPT [8], and WinCLIP [14] as baseline methods. For AnomalyGPT [8], we cast classification as a single-choice QA task by prompting the model with candidate anomaly types. For WinCLIP [14], we provide it with anomaly category descriptions and compute the similarity between patch features and anomaly category embeddings, obtaining the anomaly classification results. In the few-shot setting, we combine anomaly detection methods such as PatchCore [27], PromptAD [20], WinCLIP [14], and AnomalyCLIP [48] with the anomaly classification method ZipAF [23] as our baseline methods. We employ the threshold selection strategy proposed in AnomalyNCD [13] to determine the anomaly score threshold, and fine-tune PromptAD [20] and AnomalyCLIP [48] using few-shot anomaly samples to enhance their anomaly detection performance.
4.2 Experimental Results
Figure 5: Examples of synthetic anomaly samples generated by UniADC on the MVTec-FS dataset.
Tables 1 and 2 report the anomaly detection and classification results under zero-shot and few-shot settings, respectively. In the zero-shot setting, most baselines struggle to identify specific anomaly categories due to the absence of real anomaly samples. Although some methods incorporate Vision-Language Models (VLMs) or Large Language Models (LLMs) with anomaly priors, they still fail to achieve satisfactory anomaly detection and classification performance. In contrast, UniADC fully exploits the relevance between anomaly detection and classification, thereby achieving promising performance in both tasks. Compared to other methods, it achieves approximately a 20% improvement in classification accuracy and a 10% improvement in mIoU. Moreover, UniADC(DINO) achieves superior detection and classification accuracy over UniADC(CLIP), highlighting the remarkable performance of DINOv3 [30] in extracting fine-grained image features. In the few-shot setting, UniADC delivers substantial performance gains, even with only one anomaly sample per category, demonstrating its low cost and effective data utilization. Furthermore, it markedly outperforms other two-stage methods, showcasing the benefit of unifying anomaly detection and classification. Qualitative results in Fig. 4 further illustrate its ability to accurately localize and classify anomalous regions under various settings, emphasizing its strong generalization capability and practical applicability.
In addition, UniADC can be extended to the setting of full-shot normal samples, with the results provided in Table 3. We use the unsupervised anomaly detection methods PatchCore [27], RD4AD [4], and RealNet [46], as well as the semi-supervised method BGAD [41], in combination with the anomaly classification method ZipAF [23] as our baselines. When sufficient normal samples are available, UniADC can achieve anomaly detection performance comparable to mainstream full-shot methods. Moreover, it significantly improves anomaly classification performance, addressing the limitations of existing approaches in this aspect and demonstrating its unique practical value.
Table 3: Comparison of UniADC with alternative methods on the MVTec-FS dataset using full-shot normal samples, with $K_{a}=1$.
Method Detection Classification
I-AUC P-AUC PRO Acc mIoU
PatchCore+ZipAF [27] 98.77 98.76 95.17 69.50 39.60
RD4AD+ZipAF [4] 98.89 98.54 94.23 62.23 39.76
RealNet+ZipAF [46] 99.60 99.02 93.71 67.20 43.24
BGAD+ZipAF [41] 98.80 98.24 92.33 65.62 41.93
UniADC(CLIP) 98.85 98.73 93.76 87.01 51.59
UniADC(DINO) 99.41 98.95 94.06 87.86 51.93
Fig. 5 presents synthetic anomaly samples generated by UniADC. By incorporating an inpainting control network and reliable category consistency selection strategies, UniADC can generate high-quality anomaly samples that are mask-consistent, category-aligned, and visually diverse. These properties enable UniADC to generalize well across a wide range of anomaly detection and classification scenarios. A comparison with alternative anomaly synthesis methods is provided in the Appendix.
4.3 Ablation Study
We conduct extensive experiments to evaluate the effectiveness of the UniADC components. By default, all ablation experiments use DINOv3 [30] as the backbone of the discriminator. Table 4 investigates the impact of anomaly classification loss and category consistency selection on the performance of UniADC. When the anomaly classification loss is removed, UniADC degenerates into a standard anomaly detection model. In the zero-shot setting, image-level anomaly detection performance improves slightly, while anomaly localization performance degrades. In the few-shot setting, all metrics decline, indicating that incorporating the anomaly classification task does not compromise anomaly detection but rather enhances the robustness and accuracy of anomaly detection. Moreover, UniADC exhibits consistent performance improvement across all settings with the integration of category consistency selection, highlighting its critical role in enhancing the quality of synthesized anomaly samples. More ablation experiments and analysis can be found in the Appendix.
Table 4: Ablation results of UniADC on the MVTec-FS dataset. $\mathcal{L}_{CE}$ denotes the Cross-Entropy loss, and CCSCCS refers to the category consistency selection strategy.
Setting Component I-AUC P-AUC PRO Acc mIoU
Zero-shot $(K_{n}=2)$ w/o $\mathcal{L}_{CE}$ 97.76 96.23 87.10 - -
w/o CCSCCS 95.19 96.02 89.26 66.16 33.15
UniADC 97.09 97.04 92.15 74.74 36.66
Few-shot $(K_{n}=2,K_{a}=1)$ w/o $\mathcal{L}_{CE}$ 97.76 98.33 86.26 - -
w/o CCSCCS 96.84 98.08 88.56 79.91 45.37
UniADC 98.56 98.90 92.48 86.85 51.49
5 Conclusion
This paper integrates image anomaly detection and classification into a unified task and proposes UniADC, an innovative framework capable of anomaly synthesis, detection, and classification. UniADC consists of a training-free controllable inpainting network and a multi-task discriminator. The former can synthesize high-quality anomaly samples of specific categories guided by anomaly priors or anomaly samples, while the latter is trained on these synthesized anomaly samples, achieving unified anomaly detection and classification. Extensive experiments across multiple benchmarks validate the effectiveness and superiority of UniADC. In future work, we aim to further enhance the versatility of our approach by extending its application to diverse domains, such as medical image anomaly detection and classification.
Supplementary Material
Appendix A More Details
This section presents more experimental details, including the experimental setup for the WFDD dataset [3] and a detailed introduction to the anomaly priors provided for UniADC.
A.1 Experimental Setup for WFDD Dataset
The WFDD dataset [3] consists of four types of fabrics: Grey Cloth, Grid Cloth, Pink Flower, and Yellow Cloth. As it was originally designed for the task of anomaly detection, it lacks a detailed categorization of anomaly types. To support our proposed anomaly detection and classification task, we reclassified the anomaly categories, with examples of each anomaly category shown in Fig. S1.
A.2 Anomaly Prior
We employ three types of anomaly priors to characterize various anomaly categories, including shapes, sizes, and textual descriptions. To represent the shapes and sizes of anomalies, we define eight types of anomaly masks: Rectangle, Line, Polygon, Ellipse, Hollow Ellipse, Random Brush, Perlin Noise, and Foreground Mask. The first seven types are used to simulate various local anomalous regions, with each type defined in three sizes: Large, Medium, and Small, as illustrated in Fig. S2. Each mask size corresponds to a specific size range. Except for Perlin noise, all other mask types contain no more than two anomaly regions. Foreground Mask uses the object’s foreground as a mask to synthesize global anomalies. The textual descriptions are used to further define the appearance, attributes, and other relevant characteristics of anomalies. Table S1 presents the list of anomaly priors provided for each anomaly category. For most categories, the anomaly description is a synonym of its name or a combination of its name and the detection object, such as “Foreign matter in the bottle”. For a few anomaly categories, we introduce additional anomaly information, such as the color of the anomalous region, to aid in anomaly classification. We believe that these anomaly priors can be conveniently obtained in real-world scenarios. For example, in the industrial field, the production process and inspection experience of products can offer valuable prior knowledge for anomaly classification.
Figure S1: Examples of different anomaly categories in the WFDD dataset, with anomalous regions marked by red bounding boxes.
Figure S2: Examples of anomaly masks with different types and sizes.
Table S1: Anomaly prior list provided for each anomaly category.
Dataset Image Class Anomaly Category Description Shape Size     Dataset Image Class Anomaly Category Description Shape Size
MVTec-FS Bottle Broken large Badly broken bottle Large     MVTec-FS Screw Manipulated front Damaged screw tip
Broken small Slightly broken bottle Small     Scratch head Damaged screw head
Contamination Foreign matter in the bottle     Scratch neck Damaged screw neck
Cable Bent wire Bending copper wires     Thread side Screw thread stripping
Cable swap Cable swap Foreground     Thread top Damaged screw thread
Cut inner insulation Cracks in the internal insulation     Tile Crack Cracks on the tile
Cut outer insulation Cracks in the external insulation     Glue strip Transparent glue
Missing cable Missing cables Foreground     Gray stroke Gray-black stain Large
Missing wire Missing wires     Oil Oil stain
Poke insulation Puncture-induced holes     Rough Rough white tile Large
Capsule Crack Cracked capsule     Toothbrush Defective Missing bristles, Bending bristle, Foreign matter, A piece of thread
Faulty imprint Font misprint, Font error     Transistor Bent lead Bending metal pins
Poke A tiny hole on the capsule Ellipse Small     Cut lead Broken metal pins
Scratch Slight scratches on the capsule     Damaged case Damaged transistor
Squeeze A capsule deformed by squeezing Large     Misplaced Misplaced transistor Foreground
Carpet Color Red stain, Black stain     Wood Color Red stain, Black stain
Cut A tiny cut in carpet     Hole Hole in wood Ellipse
Hole Hole Ellipse     Liquid Transparent liquid stains
Metal contamination Metal impurities, Metal foreign matter     Scratch Scratches in wood
Thread A piece of thread Line, Hollow Ellipse Small, Medium     Zipper Broken teeth Damage to the zipper
Grid Bent Bending metal grid     Fabric border Border damage of fabric
Broken Broken defect     Fabric interior Interior damage of fabric
Glue Transparent glue     Rough Wear on the zipper
Metal contamination Metal impurities, Metal foreign matter     Split teeth Unzipped zipper
Thread A piece of thread Line, Hollow Ellipse Small, Medium     Squeezed teeth Squeezed zipper
Hazelnut Crack A hazelnut with cracks     WFDD Grey Cloth Contaminated Black stain
Cut A hazelnut with cutting marks     Flecked White spot Ellipse Small
Hole A hazelnut with a hole Ellipse     Line Black linear stain Line Small
Print A hazelnut with printed text     String Fine line Small
Leather Color Red stain, Black stain     Grid Cloth Fold Fabric with fold marks Line
Cut A tiny cut in leather     Flecked White spot Ellipse Small
Fold Fold marks in leather     String Fine line Small
Glue Transparent glue     Pink Flower Hole Hole Ellipse
Poke A puncture-induced hole Ellipse Small     Stain Small stain Small
Metal Nut Bent Cracked metal nut     Tear Zipper-like tear
Color Red stain, Black stain     Yellow cloth Fold Fabric with fold marks Line
Flip Rotating metal nut with motion blur Foreground     Stain Small stain Small
Scratch Scratches on metal nut     String Fine line Small
Pill Color Red stain, Black stain     MTD Blowhole Black hole Ellipse Small
Contamination Foreign matter on the pill     Break Surface crumpling Small
Crack Edge damage of the pill     Crack Surface cracking Small
Faulty imprint Font misprint, Font error     Fray Surface abrasion Large
Pill type Mold growth on the pill Foreground     Uneven Surface unevenness with lighting variation Large
Scratch Scratches on the pill
Appendix B More Results
This section presents additional experimental results, including comparisons with other anomaly synthesis methods, supplementary ablation studies and qualitative results, results under the open-set setting, and per-class metrics.
Table S2: Comparison of UniADC with other anomaly synthesis methods on the MVTec-FS dataset.
Setting Method Detection Classification Image Quality
I-AUC P-AUC PRO Acc mIoU IS $\uparrow$ IC-L $\uparrow$
Zero-shot $(K_{n}=2)$ CutPaste [19] 94.15 92.60 78.31 - - 1.35 0.11
NSA [29] 95.06 93.70 77.20 - - 1.50 0.14
DRÆM [42] 95.90 95.71 84.48 - - 1.53 0.16
RealNet [46] 95.83 96.11 85.76 - - 1.58 0.14
AnomalyAny [34] 93.71 93.24 82.96 52.47 23.65 1.38 0.12
UniADC 97.09 97.04 92.15 74.74 36.66 1.70 0.21
Few-shot $(K_{n}=2,K_{a}=1)$ w/o Synthesis 96.28 97.45 84.84 77.40 42.23 - -
CutPasteByMask 97.38 98.00 88.96 82.41 47.07 1.35 0.07
AnoGen [9] 97.69 98.52 88.79 85.74 50.34 1.39 0.15
DualAnoDiff [15] 96.50 98.47 89.62 85.61 50.71 1.48 0.26
AnomalyDiffusion [10] 97.16 98.21 90.16 83.07 50.82 1.50 0.21
UniADC 98.56 98.90 92.48 86.85 51.49 1.75 0.24
B.1 Comparison with Other Anomaly Synthesis Methods
Table S2 presents the performance of UniADC and other anomaly synthesis methods in terms of anomaly detection, classification, and the quality of synthetic anomaly images. For the zero-shot setting, we use CutPaste [19], NSA [29], DRÆM [42], RealNet [46], and AnomalyAny [34] as baseline methods. For the few-shot setting, we use AnoGen [9], DualAnoDiff [15], AnomalyDiffusion [10], as well as the methods without anomaly synthesis (w/o Synthesis) and CutPaste based on masks (CutPasteByMask) as our baseline methods. We use the synthetic anomaly samples from the above methods to train the multi-task discriminator and evaluate their quality using the IS and IC-LPIPS metrics, as in previous works [10, 15]. In both settings, UniADC achieves satisfactory performance and outperforms other methods in multiple aspects. Compared to w/o Synthesis and CutPasteByMask, UniADC shows significant performance improvements, demonstrating the necessity of anomaly sample augmentation and image inpainting.
Table S3: Ablation study of anomaly prior on the MVTec-FS dataset with $K_{n}=2$.
Description Shape Size I-AUC P-AUC PRO Acc mIoU
96.04 95.82 88.22 65.31 31.76
96.75 96.24 91.03 69.58 34.54
96.28 96.00 89.86 66.02 32.56
96.49 96.29 90.69 67.08 33.26
96.97 96.56 91.73 70.84 34.79
97.12 96.80 91.58 72.69 35.17
97.09 97.04 92.15 74.74 36.66
B.2 Supplementary Ablation Studies
1) Ablation Study on Anomaly Prior. Table S3 investigates the impact of different types of anomaly priors on the performance of UniADC. When anomaly descriptions are unavailable, we use simple combinations of anomaly category names and detection objects as prompts. The experimental results show that anomaly descriptions have the most significant impact on the performance of UniADC, especially in anomaly classification. The introduction of anomaly descriptions enables UniADC to achieve improvements of 4.27% in classification accuracy and 2.78% in mIoU. Compared to anomaly descriptions, shape and size priors are generally difficult to apply to certain types of anomalies. For example, it is hard to predetermine the shapes and sizes of anomaly types such as “Scratch” and “Crack”. Moreover, in our experiments, the size prior yields greater performance improvements for UniADC than the shape prior. When all three types of anomaly priors are applied, UniADC achieves the best overall performance.
Table S4: Ablation study of the noise factor $\gamma$ on the MVTec-FS dataset.
Setting $\gamma$ I-AUC P-AUC PRO Acc mIoU
Zero-shot $(K_{n}=2)$ $(0,0.2]$ 93.19 95.76 85.43 61.46 30.89
$(0.2,0.4]$ 96.35 96.78 91.03 73.67 36.84
$(0.4,0.6]$ 97.09 97.04 92.15 74.74 36.66
$(0.6,0.8]$ 97.14 96.65 91.58 74.40 35.81
$(0.8,1.0]$ 96.83 96.38 90.38 72.91 34.83
Few-shot $(K_{n}=2,K_{a}=1)$ $(0,0.2]$ 97.85 98.53 90.86 84.44 49.34
$(0.2,0.4]$ 98.69 98.71 91.92 85.89 50.85
$(0.4,0.6]$ 98.56 98.90 92.48 86.85 51.49
$(0.6,0.8]$ 98.42 98.84 92.18 86.54 51.55
$(0.8,1.0]$ 98.21 98.77 88.95 85.03 50.40
Figure S3: Examples of synthetic anomaly images generated with different noise factors $\gamma$.
2) Ablation Study on Noise Factor. Table S4 presents the impact of the noise factor $\gamma$ on the performance of UniADC. A larger $\gamma$ increases the diversity of synthetic anomaly samples while also causing greater deviation from the original image distribution, as shown in Fig. S3, which may undermine UniADC’s ability to detect hard anomaly cases. Conversely, a smaller $\gamma$ may result in synthetic anomaly samples that do not align well with the provided anomaly description, leading to false positive anomalies or limited diversity. In our experiments, we uniformly sample $\gamma$ between 0.4 and 0.6 to ensure a balance between quality and diversity.
Table S5: Ablation study of the mini-batch size $B$ on the MVTec-FS dataset.
Setting $\qquad B\qquad$ I-AUC P-AUC PRO Acc mIoU
Zero-shot $(K_{n}=2)$ 8 95.93 96.55 90.55 71.90 33.17
16 96.29 96.80 90.38 73.12 35.14
32 97.09 97.04 92.15 74.74 36.66
64 97.07 96.98 92.47 74.91 36.98
Few-shot $(K_{n}=2,K_{a}=1)$ 8 97.97 98.67 90.20 84.72 50.17
16 98.04 98.72 90.49 85.94 50.63
32 98.56 98.90 92.48 86.85 51.49
64 99.02 98.70 92.81 86.93 51.12
Table S6: Ablation study of the number of synthetic anomaly samples on the MVTec-FS dataset.
Setting Number of Synthetic Samples I-AUC P-AUC PRO Acc mIoU
Zero-shot $(K_{n}=2)$ 2 93.25 95.89 86.10 64.19 30.37
4 95.99 96.13 88.75 69.63 33.20
8 96.34 96.85 91.73 73.55 35.79
16 97.09 97.04 92.15 74.74 36.66
32 97.16 96.93 92.04 74.82 36.86
Few-shot $(K_{n}=2,K_{a}=1)$ 2 97.25 98.05 87.90 83.05 46.70
4 98.09 98.05 91.02 84.94 49.34
8 98.27 98.16 91.53 85.59 51.03
16 98.56 98.90 92.48 86.85 51.49
32 98.72 98.75 92.20 86.94 51.80
3) Ablation Study on Mini-batch Size and the Number of Synthetic Samples. Tables S5 and S6 respectively present the ablation study results on the mini-batch size in category consistency selection and the number of synthetic anomaly samples. Increasing the mini-batch size and the number of synthetic samples can improve the quality and diversity of training data, thereby enhancing the performance of UniADC. However, this also significantly increases the training cost. To ensure the performance of UniADC while reducing computational cost, we set the mini-batch size and the number of synthetic anomaly samples to 32 and 16, respectively. We implement UniADC using PyTorch, and it takes approximately 35 seconds to generate an anomaly image on a single Nvidia RTX 4090 GPU. The average training time of UniADC (including anomaly sample synthesis and discriminator training) is about 1 GPU hour. During the inference stage, the main computational overhead of UniADC lies in the forward pass of the vision-language backbone, and its inference speed is comparable to that of other VLM-based methods, such as AnomalyCLIP [48].
4) Ablation Study on Anomaly Score Threshold. Fig. S4 illustrates the effect of varying the anomaly score threshold on the anomaly classification performance of UniADC. Compared to other methods [14, 20], UniADC demonstrates improved robustness to threshold variations, consistently achieving reliable performance when the threshold is set between 0.4 and 0.6, which simplifies the threshold selection process and improves practical feasibility. In our experiments, we simply fixed the threshold at 0.5.
Figure S4: Ablation study of the anomaly score threshold on the MVTec-FS dataset.
B.3 Open-set Anomaly Detection and Classification
For open-set anomaly detection, a model is trained on only a subset of anomaly types and is expected to generalize to detect unseen anomalies [6]. We extend this setting to anomaly detection and classification, where a model is designed to effectively detect both seen and unseen anomalies while ensuring correct classification of the seen anomalies. We propose a simple yet effective approach to adapt UniADC for open-set anomaly detection and classification tasks. Specifically, during training, we use unconstrained anomaly synthesis methods (e.g., DRÆM [42]) to generate class-agnostic anomalous images and align them with an additional learnable text embedding. During inference, the learnable text embedding is used to match all unseen anomaly types and categorize them as “Other”.
We conduct experiments on the Hazelnut class of the MVTec-FS dataset [23] to validate the effectiveness of our proposed method. We use Crack, Cut, and Hole as seen anomalies, and Print as the unseen anomaly. We adopt the zero-shot setting ($K_{n}=2$), where no prior knowledge about the Print category is provided during training. Tab. S7 presents the anomaly detection and classification results of UniADC under both open-set and closed-set settings. In the open-set setting, UniADC achieves anomaly detection and classification performance comparable to the closed-set, demonstrating that it generalizes well to unseen anomalies while avoiding their interference with anomaly classification. Fig. S5 presents qualitative results of UniADC, showing that it ensures accurate detection and classification of seen anomalies, while aligning unseen anomalies with the learnable text embedding to identify them as the “Other” type.
Table S7: Comparison of UniADC performance on the Hazelnut class of MVTec-FS dataset under open- and closed-set settings.
Setting I-AUC P-AUC PRO Acc (seen)
Normal Crack Cut Hole AVG
Closed-set 99.63 97.53 91.29 97.50 55.56 62.50 88.89 86.36
Open-set 99.41 99.09 95.21 95.00 44.44 62.50 88.89 83.33
Figure S5: Qualitative results of UniADC under the setting of open-set anomaly detection and classification.
B.4 Supplementary Qualitative Results
Figs. S6S8 provide the qualitative results of UniADC on the MVTec-FS [23], MTD [12], and WFDD [3] datasets, while Fig. S9 shows some examples of synthetic anomaly samples. These qualitative results consistently demonstrate the outstanding performance of UniADC in anomaly detection, classification, and synthesis, further highlighting its robustness and broad applicability across diverse anomaly detection and classification tasks.
B.5 Metrics for Each Class
Tables S8S39 present the class-specific anomaly detection and classification metrics of UniADC on the MVTec-FS and WFDD datasets. UniADC achieves satisfactory performance on most classes, providing a valuable reference for anomaly detection and root cause analysis in real-world scenarios.
Figure S6: Qualitative comparison of UniADC and other methods on the Hazelnut, Pill, and Metal Nut classes of the MVTec-FS dataset.
Figure S7: Qualitative comparison of UniADC and other methods on the Screw, Leather, and Wood classes of the MVTec-FS dataset.
Figure S8: Qualitative comparison of UniADC and other methods on the MTD and WFDD datasets.
Figure S9: Examples of synthetic anomaly samples generated by UniADC on the MVTec-FS, MTD, and WFDD datasets.
Table S8: Anomaly detection and classification results of UniADC(CLIP) on the MVTec-FS dataset under the zero-shot $(K_{n}=1)$ setting.
I-AUC P-AUC PRO Acc mIoU
Bottle 99.52 93.85 85.24 72.55 42.41
Cable 93.67 94.53 84.56 62.64 33.59
Capsule 90.98 97.12 92.80 43.42 24.16
Carpet 99.57 99.09 97.74 70.00 30.37
Grid 99.12 98.73 94.31 62.50 24.53
Hazelnut 99.93 98.42 95.85 75.67 29.92
Leather 100.0 99.85 98.86 77.63 39.34
Metal Nut 99.60 96.69 78.57 79.10 40.59
Pill 98.30 98.40 96.27 56.98 28.41
Screw 77.12 97.68 90.90 35.35 17.92
Tile 99.63 99.17 94.58 82.43 39.87
Toothbrush 95.00 97.05 69.96 92.59 62.76
Transistor 80.33 78.81 70.79 60.00 21.26
Wood 97.37 97.92 88.78 73.81 40.89
Zipper 95.37 97.63 92.08 46.34 17.52
AVG 95.03 96.33 88.75 66.07 32.90
Table S9: Anomaly detection and classification results of UniADC(CLIP) on the MVTec-FS dataset under the zero-shot $(K_{n}=2)$ setting.
I-AUC P-AUC PRO Acc mIoU
Bottle 98.06 96.54 84.09 76.47 39.18
Cable 93.06 94.98 86.98 66.33 27.81
Capsule 89.01 94.83 90.08 61.84 26.84
Carpet 100.0 98.74 96.86 71.43 29.19
Grid 99.82 98.71 95.49 70.83 30.34
Hazelnut 99.56 99.18 94.88 85.14 44.63
Leather 100.0 99.87 97.17 75.00 35.64
Metal Nut 99.70 97.55 84.20 83.58 41.15
Pill 96.83 98.25 97.20 65.12 26.20
Screw 80.91 97.48 88.24 45.45 20.81
Tile 100.0 99.16 94.53 81.08 48.93
Toothbrush 95.56 98.16 67.10 85.19 63.61
Transistor 79.75 77.45 60.85 75.00 22.31
Wood 98.12 98.84 95.82 80.95 50.89
Zipper 98.65 98.23 94.36 51.22 19.24
AVG 95.27 96.53 88.52 71.64 35.12
Table S10: Anomaly detection and classification results of UniADC(CLIP) on the MVTec-FS dataset under the zero-shot $(K_{n}=4)$ setting.
I-AUC P-AUC PRO Acc mIoU
Bottle 99.20 94.69 87.26 74.51 40.20
Cable 91.99 93.31 82.95 74.49 30.52
Capsule 96.23 97.55 94.14 63.16 28.08
Carpet 100.0 98.82 97.59 72.86 30.27
Grid 98.41 99.04 95.48 68.75 23.19
Hazelnut 99.85 98.79 95.48 81.08 39.17
Leather 99.15 99.65 99.05 76.32 39.82
Metal Nut 97.79 96.22 90.50 82.09 40.20
Pill 98.64 98.93 95.62 63.95 26.92
Screw 86.42 98.57 91.65 49.49 20.35
Tile 100.0 98.60 95.79 85.14 45.41
Toothbrush 97.78 95.18 75.49 92.59 62.41
Transistor 83.50 80.84 78.72 75.00 22.92
Wood 98.50 98.88 95.30 88.10 57.81
Zipper 95.26 97.93 92.59 52.44 19.76
AVG 96.18 96.47 91.17 73.33 35.14
Table S11: Anomaly detection and classification results of UniADC(CLIP) on the MVTec-FS dataset under the few-shot $(K_{n}=1,K_{a}=1)$ setting.
I-AUC P-AUC PRO Acc mIoU
Bottle 99.03 99.53 90.68 84.31 61.97
Cable 98.89 96.78 90.21 93.88 52.38
Capsule 97.70 98.24 94.09 71.05 32.34
Carpet 100.0 99.66 91.38 90.00 48.08
Grid 99.82 99.46 97.57 85.42 36.86
Hazelnut 100.0 98.50 88.62 100.0 55.76
Leather 100.0 99.85 95.80 88.16 53.67
Metal Nut 100.0 99.75 90.97 88.06 65.96
Pill 95.76 99.62 84.66 65.12 38.28
Screw 83.77 97.79 87.61 46.46 23.47
Tile 100.0 99.42 93.79 100.0 77.38
Toothbrush 95.00 97.10 71.59 92.59 59.24
Transistor 96.50 93.75 74.14 87.50 24.65
Wood 99.62 98.99 96.54 92.86 51.61
Zipper 99.30 99.54 98.14 85.37 47.66
AVG 97.69 98.53 89.72 84.72 48.62
Table S12: Anomaly detection and classification results of UniADC(CLIP) on the MVTec-FS dataset under the few-shot $(K_{n}=2,K_{a}=1)$ setting.
I-AUC P-AUC PRO Acc mIoU
Bottle 99.35 99.51 84.85 84.31 63.35
Cable 98.08 97.06 86.59 94.90 51.09
Capsule 98.11 98.54 95.52 65.79 35.10
Carpet 100.0 99.50 98.93 87.14 46.97
Grid 100.0 99.52 94.52 85.42 34.83
Hazelnut 100.0 98.16 87.54 98.65 57.38
Leather 100.0 99.85 97.10 88.16 49.87
Metal Nut 100.0 99.67 87.34 89.55 63.53
Pill 97.45 99.46 91.53 62.79 38.66
Screw 87.89 98.40 86.62 60.61 31.29
Tile 100.0 99.45 93.00 100.0 78.11
Toothbrush 98.33 96.91 85.50 92.59 64.82
Transistor 96.50 94.70 83.74 86.25 27.10
Wood 99.81 99.18 95.58 92.86 49.89
Zipper 99.73 99.54 91.30 89.02 41.97
AVG 98.35 98.63 90.64 85.20 48.93
Table S13: Anomaly detection and classification results of UniADC(CLIP) on the MVTec-FS dataset under the few-shot $(K_{n}=2,K_{a}=2)$ setting.
I-AUC P-AUC PRO Acc mIoU
Bottle 99.03 99.49 91.75 86.27 60.96
Cable 98.39 96.76 93.63 93.88 51.90
Capsule 97.29 98.83 95.44 78.95 36.03
Carpet 100.0 99.50 95.08 92.86 49.76
Grid 99.82 99.39 97.45 87.50 38.92
Hazelnut 100.0 99.71 95.57 100.0 75.10
Leather 100.0 99.84 98.46 93.42 58.77
Metal Nut 100.0 99.75 94.96 88.06 65.27
Pill 99.04 99.81 90.94 77.91 49.99
Screw 95.04 99.13 95.06 72.73 36.34
Tile 100.0 99.41 94.54 100.0 79.37
Toothbrush 96.11 97.39 78.95 92.59 65.9
Transistor 97.50 95.51 85.46 92.50 36.52
Wood 99.62 99.29 97.20 97.62 63.92
Zipper 99.35 99.42 95.10 93.90 55.24
AVG 98.75 98.88 93.31 89.88 54.93
Table S14: Anomaly detection and classification results of UniADC(CLIP) on the MVTec-FS dataset under the few-shot $(K_{n}=4,K_{a}=1)$ setting.
I-AUC P-AUC PRO Acc mIoU
Bottle 100.0 99.55 93.76 82.35 65.97
Cable 99.04 97.00 89.19 94.90 53.99
Capsule 98.85 98.23 93.57 68.42 33.75
Carpet 100.0 99.67 98.75 88.57 46.52
Grid 99.29 99.49 97.22 87.50 40.92
Hazelnut 100.0 98.27 95.95 97.30 55.50
Leather 100.0 99.85 95.94 86.84 54.37
Metal Nut 100.0 99.72 90.39 88.06 64.02
Pill 98.08 99.62 89.39 70.93 41.00
Screw 92.60 98.89 82.67 55.56 26.84
Tile 100.0 99.50 92.37 100.0 76.21
Toothbrush 96.67 97.76 78.30 92.59 65.16
Transistor 94.50 92.34 82.60 87.50 28.71
Wood 100.0 98.97 96.22 95.24 53.34
Zipper 99.52 99.46 95.03 87.80 43.81
AVG 98.57 98.55 91.42 85.57 50.01
Table S15: Anomaly detection and classification results of UniADC(CLIP) on the MVTec-FS dataset with full-shot normal samples and $K_{a}=1$.
I-AUC P-AUC PRO Acc mIoU
Bottle 100.0 99.68 92.40 82.35 65.17
Cable 99.35 96.77 94.87 94.90 55.04
Capsule 99.51 99.00 97.08 71.05 34.34
Carpet 100.0 99.34 95.58 88.57 47.77
Grid 99.65 99.39 96.92 89.58 42.56
Hazelnut 100.0 98.02 92.78 98.65 57.54
Leather 100.0 99.83 97.24 88.16 52.82
Metal Nut 100.0 99.74 94.11 94.03 66.48
Pill 99.04 99.68 92.54 72.09 45.28
Screw 91.34 98.92 90.67 62.63 30.26
Tile 100.0 99.44 96.61 100.0 76.79
Toothbrush 96.11 98.81 83.80 92.59 66.97
Transistor 98.25 94.12 87.52 90.00 28.23
Wood 100.0 98.65 96.34 95.24 58.06
Zipper 99.46 99.54 98.00 85.37 46.49
AVG 98.85 98.73 93.76 87.01 51.59
Table S16: Anomaly detection and classification results of UniADC(DINO) on the MVTec-FS dataset under the zero-shot $(K_{n}=1)$ setting.
I-AUC P-AUC PRO Acc mIoU
Bottle 99.03 93.73 80.66 80.39 47.82
Cable 94.29 94.46 88.32 55.10 22.53
Capsule 94.26 98.02 93.16 59.21 26.57
Carpet 100.0 99.29 98.43 85.71 39.34
Grid 100.0 99.61 98.45 62.50 25.09
Hazelnut 99.19 98.24 80.86 75.68 38.12
Leather 100.0 99.80 96.53 80.26 39.12
Metal Nut 94.65 97.16 96.34 55.22 37.30
Pill 96.83 92.32 95.57 50.00 20.50
Screw 81.96 97.51 87.82 31.31 20.34
Tile 99.63 97.83 90.58 78.38 35.43
Toothbrush 95.00 98.87 64.09 92.59 70.73
Transistor 92.75 77.29 72.72 81.25 24.50
Wood 98.31 98.38 97.27 85.71 60.13
Zipper 99.68 99.09 96.53 51.22 18.35
AVG 96.37 96.11 89.16 68.30 35.06
Table S17: Anomaly detection and classification results of UniADC(DINO) on the MVTec-FS dataset under the zero-shot $(K_{n}=2)$ setting.
I-AUC P-AUC PRO Acc mIoU
Bottle 98.39 97.81 93.78 82.35 54.97
Cable 92.53 96.26 87.31 65.31 23.65
Capsule 98.52 98.78 97.59 61.84 32.78
Carpet 99.66 98.71 97.98 81.43 34.54
Grid 100.0 99.61 96.57 72.92 31.05
Hazelnut 99.63 97.53 91.29 85.14 38.84
Leather 99.93 99.52 98.39 82.89 41.56
Metal Nut 99.39 98.05 97.08 79.10 42.24
Pill 95.98 88.81 96.12 58.14 23.38
Screw 88.98 98.25 91.98 55.56 26.93
Tile 99.48 98.66 94.62 78.38 34.05
Toothbrush 99.44 97.19 79.21 96.30 68.10
Transistor 92.92 89.11 69.53 77.50 21.62
Wood 93.61 98.95 96.26 88.10 56.52
Zipper 97.95 98.36 94.50 56.10 19.73
AVG 97.09 97.04 92.15 74.74 36.66
Table S18: Anomaly detection and classification results of UniADC(DINO) on the MVTec-FS dataset under the zero-shot $(K_{n}=4)$ setting.
I-AUC P-AUC PRO Acc mIoU
Bottle 99.35 96.82 94.11 82.35 56.14
Cable 92.07 95.40 90.47 71.43 29.28
Capsule 97.95 97.00 93.41 63.16 31.50
Carpet 100.0 99.63 95.21 72.86 30.93
Grid 100.0 99.60 97.57 66.67 28.54
Hazelnut 99.56 98.62 95.24 85.14 36.31
Leather 100.0 99.78 99.63 90.78 49.25
Metal Nut 99.90 98.36 94.95 82.09 37.84
Pill 97.79 96.11 97.90 56.98 21.66
Screw 88.44 97.86 91.35 55.56 20.66
Tile 98.45 98.07 95.64 86.49 43.62
Toothbrush 100.0 96.90 73.31 100.0 63.42
Transistor 95.00 88.84 79.31 86.25 26.01
Wood 98.12 99.16 97.04 88.10 59.55
Zipper 98.06 98.27 95.05 57.32 23.80
AVG 97.65 97.36 92.68 76.35 37.23
Table S19: Anomaly detection and classification results of UniADC(DINO) on the MVTec-FS dataset under the few-shot $(K_{n}=1,K_{a}=1)$ setting.
I-AUC P-AUC PRO Acc mIoU
Bottle 99.52 99.71 88.05 82.35 61.95
Cable 96.74 96.77 90.09 92.86 54.27
Capsule 99.51 99.26 96.03 76.32 36.23
Carpet 100.0 99.81 98.04 81.43 45.19
Grid 100.0 99.51 96.77 91.67 41.38
Hazelnut 100.0 98.36 93.94 100.0 54.67
Leather 100.0 99.84 96.65 93.42 50.08
Metal Nut 99.90 99.74 94.22 88.06 65.78
Pill 95.48 99.67 79.47 66.28 40.39
Screw 85.58 99.17 88.31 54.55 30.12
Tile 100.0 99.44 96.31 98.65 74.52
Toothbrush 100.0 97.73 81.59 100.0 67.76
Transistor 100.0 96.68 92.18 91.25 30.78
Wood 99.62 99.03 94.94 92.86 57.39
Zipper 99.95 99.72 97.24 91.46 58.72
AVG 98.42 98.96 92.26 86.74 51.28
Table S20: Anomaly detection and classification results of UniADC(DINO) on the MVTec-FS dataset under the few-shot $(K_{n}=2,K_{a}=1)$ setting.
I-AUC P-AUC PRO Acc mIoU
Bottle 100.0 99.69 90.82 82.35 61.38
Cable 97.28 97.22 85.47 93.88 56.36
Capsule 98.28 99.18 95.61 75.00 38.77
Carpet 100.0 99.41 98.48 81.43 43.57
Grid 100.0 99.58 96.91 91.67 37.38
Hazelnut 100.0 99.12 95.51 98.65 54.04
Leather 100.0 99.83 96.21 93.42 50.59
Metal Nut 100.0 99.77 93.91 89.55 65.82
Pill 97.17 99.60 87.16 68.60 40.80
Screw 88.10 98.98 86.88 62.63 31.49
Tile 100.0 99.31 96.79 98.65 75.22
Toothbrush 98.33 97.13 87.94 92.59 68.74
Transistor 99.92 96.25 86.5 91.25 29.50
Wood 99.25 98.73 94.72 92.86 57.74
Zipper 100.0 99.72 94.34 90.24 60.95
AVG 98.56 98.90 92.48 86.85 51.49
Table S21: Anomaly detection and classification results of UniADC(DINO) on the MVTec-FS dataset under the few-shot $(K_{n}=2,K_{a}=2)$ setting.
I-AUC P-AUC PRO Acc mIoU
Bottle 100.0 99.67 94.18 84.31 55.23
Cable 98.05 97.93 89.26 87.76 45.81
Capsule 98.93 99.10 92.47 77.63 40.49
Carpet 99.66 99.70 98.77 88.57 50.76
Grid 100.0 99.63 96.01 87.50 41.06
Hazelnut 100.0 99.78 95.58 97.30 70.42
Leather 100.0 99.82 96.77 93.42 52.75
Metal Nut 100.0 99.76 95.22 88.06 66.13
Pill 99.21 99.81 93.64 73.26 51.18
Screw 89.96 99.13 95.36 72.73 36.34
Tile 100.0 99.30 98.23 100.0 77.04
Toothbrush 100.0 97.30 83.60 96.30 66.82
Transistor 99.92 96.76 85.19 91.25 38.44
Wood 100.0 99.12 94.82 97.62 60.69
Zipper 100.0 99.67 94.12 95.12 59.93
AVG 99.05 99.10 93.55 88.72 54.21
Table S22: Anomaly detection and classification results of UniADC(DINO) on the MVTec-FS dataset under the few-shot $(K_{n}=4,K_{a}=1)$ setting.
I-AUC P-AUC PRO Acc mIoU
Bottle 99.68 99.65 92.64 82.35 64.09
Cable 97.82 96.17 86.18 92.86 56.49
Capsule 99.10 99.24 92.04 76.32 39.26
Carpet 100.0 99.23 98.07 84.29 45.52
Grid 100.0 99.59 96.08 91.67 38.76
Hazelnut 100.0 99.03 98.33 95.95 55.39
Leather 100.0 99.83 96.81 89.47 54.12
Metal Nut 100.0 99.79 92.50 86.57 65.63
Pill 97.51 99.66 92.65 69.77 43.59
Screw 87.55 98.97 89.40 57.58 27.68
Tile 100.0 98.55 97.52 100.0 74.99
Toothbrush 100.0 98.26 86.35 100.0 69.02
Transistor 99.83 97.00 89.72 91.25 33.87
Wood 99.06 98.97 93.59 92.86 58.14
Zipper 100.0 99.74 98.44 91.46 59.02
AVG 98.70 98.91 93.35 86.83 52.37
Table S23: Anomaly detection and classification results of UniADC(DINO) on the MVTec-FS dataset with full-shot normal samples and $K_{a}=1$.
I-AUC P-AUC PRO Acc mIoU
Bottle 100.0 99.71 94.61 78.43 58.28
Cable 97.66 96.48 90.91 94.90 56.25
Capsule 99.26 99.33 95.31 75.00 42.60
Carpet 99.92 99.23 96.55 82.86 42.77
Grid 100.0 99.59 96.71 85.42 42.41
Hazelnut 100.0 98.65 92.32 97.30 55.91
Leather 100.0 99.85 96.04 92.11 52.04
Metal Nut 100.0 99.78 94.00 89.55 66.52
Pill 99.32 99.62 89.46 72.09 44.18
Screw 95.21 99.00 90.42 74.75 31.86
Tile 100.0 99.34 97.99 100.0 75.35
Toothbrush 100.0 97.10 94.60 100.0 67.21
Transistor 100.0 97.87 89.15 90.00 34.11
Wood 99.81 98.99 97.04 95.24 53.83
Zipper 100.0 99.68 95.79 90.24 55.70
AVG 99.41 98.95 94.06 87.86 51.93
Table S24: Anomaly detection and classification results of UniADC(CLIP) on the WFDD dataset under the zero-shot $(K_{n}=1)$ setting.
I-AUC P-AUC PRO Acc mIoU
Grey Cloth 99.89 99.08 85.75 81.25 38.05
Grid Cloth 94.51 98.10 78.98 90.00 44.08
Pink Flower 95.18 99.90 90.85 85.71 55.14
Yellow Cloth 98.46 98.92 92.54 89.82 44.93
AVG 97.01 99.00 87.03 86.70 45.55
Table S25: Anomaly detection and classification results of UniADC(CLIP) on the WFDD dataset under the zero-shot $(K_{n}=2)$ setting.
I-AUC P-AUC PRO Acc mIoU
Grey Cloth 99.11 98.67 84.78 92.19 42.70
Grid Cloth 95.43 98.56 79.80 86.00 41.23
Pink Flower 97.27 99.90 91.49 89.29 62.02
Yellow Cloth 98.39 98.79 92.47 90.42 46.64
AVG 97.55 98.98 87.14 89.48 48.15
Table S26: Anomaly detection and classification results of UniADC(CLIP) on the WFDD dataset under the zero-shot $(K_{n}=4)$ setting.
I-AUC P-AUC PRO Acc mIoU
Grey Cloth 99.89 98.95 90.16 95.32 42.74
Grid Cloth 94.42 98.51 75.61 83.00 41.52
Pink Flower 98.96 99.67 96.86 92.86 62.84
Yellow Cloth 99.34 99.22 90.51 92.22 45.85
AVG 98.15 99.09 88.29 90.85 48.24
Table S27: Anomaly detection and classification results of UniADC(CLIP) on the WFDD dataset under the few-shot $(K_{n}=1,K_{a}=1)$ setting.
I-AUC P-AUC PRO Acc mIoU
Grey Cloth 99.78 98.08 88.76 92.19 39.84
Grid Cloth 95.76 98.68 92.87 89.00 40.55
Pink Flower 98.70 99.97 95.56 96.43 68.70
Yellow Cloth 99.77 98.60 94.63 96.41 45.93
AVG 98.50 98.83 92.96 93.51 48.76
Table S28: Anomaly detection and classification results of UniADC(CLIP) on the WFDD dataset under the few-shot $(K_{n}=2,K_{a}=1)$ setting.
I-AUC P-AUC PRO Acc mIoU
Grey Cloth 99.67 98.80 95.53 92.19 42.92
Grid Cloth 97.41 98.62 94.95 89.00 41.42
Pink Flower 98.96 99.98 93.74 96.43 68.46
Yellow Cloth 99.85 99.23 94.35 97.01 46.04
AVG 98.97 99.16 94.64 93.66 49.71
Table S29: Anomaly detection and classification results of UniADC(CLIP) on the WFDD dataset under the few-shot $(K_{n}=2,K_{a}=2)$ setting.
I-AUC P-AUC PRO Acc mIoU
Grey Cloth 99.11 98.89 95.70 93.75 42.85
Grid Cloth 98.75 99.11 93.07 90.00 43.64
Pink Flower 98.70 99.97 95.06 96.43 67.73
Yellow Cloth 100.0 99.40 95.55 99.40 50.02
AVG 99.14 99.34 94.85 94.90 51.06
Table S30: Anomaly detection and classification results of UniADC(CLIP) on the WFDD dataset under the few-shot $(K_{n}=4,K_{a}=1)$ setting.
I-AUC P-AUC PRO Acc mIoU
Grey Cloth 99.78 98.67 92.77 92.19 42.24
Grid Cloth 98.06 98.89 94.81 90.00 42.79
Pink Flower 99.61 99.97 95.55 96.43 67.80
Yellow Cloth 99.94 99.24 94.80 97.01 46.92
AVG 99.35 99.19 94.48 93.91 49.94
Table S31: Anomaly detection and classification results of UniADC(CLIP) on the WFDD dataset with full-shot normal samples and $K_{a}=1$.
I-AUC P-AUC PRO Acc mIoU
Grey Cloth 99.67 98.91 94.10 92.19 42.21
Grid Cloth 98.30 98.55 95.79 90.00 44.99
Pink Flower 99.87 99.98 95.34 98.21 67.60
Yellow Cloth 99.94 99.22 96.88 98.20 47.46
AVG 99.45 99.17 95.53 94.65 50.57
Table S32: Anomaly detection and classification results of UniADC(DINO) on the WFDD dataset under the zero-shot $(K_{n}=1)$ setting.
I-AUC P-AUC PRO Acc mIoU
Grey Cloth 99.11 99.11 85.80 71.88 36.15
Grid Cloth 95.76 99.29 92.48 92.00 52.53
Pink Flower 97.66 99.94 93.57 94.64 66.77
Yellow Cloth 99.77 99.49 96.65 97.01 53.24
AVG 98.08 99.46 92.13 88.88 52.17
Table S33: Anomaly detection and classification results of UniADC(DINO) on the WFDD dataset under the zero-shot $(K_{n}=2)$ setting.
I-AUC P-AUC PRO Acc mIoU
Grey Cloth 99.89 99.22 84.57 73.44 38.96
Grid Cloth 94.91 99.15 93.18 90.00 51.90
Pink Flower 98.18 99.91 94.48 96.43 69.34
Yellow Cloth 99.84 99.57 97.15 97.01 53.33
AVG 98.21 99.46 92.35 89.22 53.38
Table S34: Anomaly detection and classification results of UniADC(DINO) on the WFDD dataset under the zero-shot $(K_{n}=4)$ setting.
I-AUC P-AUC PRO Acc mIoU
Grey Cloth 100.0 99.33 90.00 73.44 40.04
Grid Cloth 94.38 98.72 95.09 92.00 51.85
Pink Flower 100.0 99.96 95.67 100.0 74.90
Yellow Cloth 100.0 99.56 98.26 98.80 54.98
AVG 98.60 99.39 94.76 91.06 55.44
Table S35: Anomaly detection and classification results of UniADC(DINO) on the WFDD dataset under the few-shot $(K_{n}=1,K_{a}=1)$ setting.
I-AUC P-AUC PRO Acc mIoU
Grey Cloth 99.67 98.74 89.51 92.19 42.52
Grid Cloth 99.72 99.41 96.11 94.00 45.38
Pink Flower 100.0 99.98 95.19 100.0 68.18
Yellow Cloth 99.99 99.35 95.48 98.20 46.03
AVG 99.85 99.37 94.07 96.10 50.53
Table S36: Anomaly detection and classification results of UniADC(DINO) on the WFDD dataset under the few-shot $(K_{n}=2,K_{a}=1)$ setting.
I-AUC P-AUC PRO Acc mIoU
Grey Cloth 99.78 99.03 90.87 96.88 42.32
Grid Cloth 99.72 99.45 97.17 95.00 49.43
Pink Flower 100.0 99.97 92.74 98.21 68.35
Yellow Cloth 99.97 99.46 95.68 98.80 47.02
AVG 99.87 99.48 94.12 97.22 51.78
Table S37: Anomaly detection and classification results of UniADC(DINO) on the WFDD dataset under the few-shot $(K_{n}=2,K_{a}=2)$ setting.
I-AUC P-AUC PRO Acc mIoU
Grey Cloth 99.89 99.00 94.57 98.44 46.38
Grid Cloth 99.84 99.44 96.82 95.00 46.76
Pink Flower 100.0 99.98 91.60 100.0 67.73
Yellow Cloth 100.0 99.54 93.77 97.60 46.78
AVG 99.93 99.49 94.19 97.76 51.91
Table S38: Anomaly detection and classification results of UniADC(DINO) on the WFDD dataset under the few-shot $(K_{n}=4,K_{a}=1)$ setting.
I-AUC P-AUC PRO Acc mIoU
Grey Cloth 99.78 98.91 94.05 96.88 43.22
Grid Cloth 99.68 99.53 94.24 95.00 46.77
Pink Flower 100.0 99.97 94.17 100.0 69.40
Yellow Cloth 100.0 99.38 95.88 98.80 47.79
AVG 99.87 99.45 94.59 97.67 51.80
Table S39: Anomaly detection and classification results of UniADC(DINO) on the WFDD dataset with full-shot normal samples and $K_{a}=1$.
I-AUC P-AUC PRO Acc mIoU
Grey Cloth 100.0 98.82 95.95 97.14 40.78
Grid Cloth 99.96 99.51 94.90 95.00 48.75
Pink Flower 100.0 99.97 95.92 100.0 70.59
Yellow Cloth 99.97 99.44 94.58 98.80 49.85
AVG 99.98 99.44 95.34 97.74 52.49